The probability that a salesman makes a sale on a visit to a prospective client is 0.2. What is
the probability that, in 2 visits, of:
i) making no sales
ii)making more than one sale
1
Expert's answer
2020-10-18T17:40:37-0400

solution


Probability of success "p=0.2"


Let "X" be the number of sales made out of the 2 attempts.



"P(X=x) =\ _nC_x * p^x *(1-p)^{n-x}"

Part i) making no sales


"P(X=0)"


"=\ _2C_0 * 0.2^0 *(0.8)^{2}= 0.64"

answer: 0.64


Part ii) making more than 1 sale


"P(X>1) = P(X=2)"


"=\ _2C_2 * 0.2^2 *(0.8)^{0}= 0.04"


answer: 0.04